# Initial Guessing Bias: How Untrained Networks Favor Some Classes

### v4

@@ 1 Introduction
Figure 1: Initial Guessing Bias (IGB). Consider a task where we classify a binary dataset using an untrained network. Does it assign half of the examples to each class, or does it privilege one class? The answer depends on the model design. In the top-left, we classify a binary dataset with an untrained network without IGB. This model will generally assign half of the examples to each class (histogram on the top-center). In the bottom-left, we classify the same dataset using an untrained network with IGB. In this case, most of the guesses will usually go to one of the two classes (histogram on the bottom-center). As an example, we take the dog/cat classes (label 0 / label 11) from CIFAR10, and pass them through an untrained CNN with 2 layers, each followed by pooling. The non-IGB model uses $\tanh$ activations and average pooling, the IGB model uses ReLU and max pooling. We show in the center-right the distribution over different initializations, $p_{f_{0}}\left({x}\right)$, of the fraction $f_{0}$ of times that each model guessed dog (equivalently, $f_{1}=1-f_{0}$ indicates the fraction of images guessed as cat). While for the non-IGB models, $f_{0}$ is most often 50%, with IGB it most often is either 0% or 100%.
@@ 3 Preliminaries
Figure 2: Left: IGB and Performance Bound: Diagram of the accessible performance range conditioned on the behavior of $\max_{c}\{f_{c}\left(\mathcal{W}_{t}\right)\}$ in a balanced binary dataset, in accordance with Eq. (1). Right: Comparison of the trend of $\max_{c}\{f_{c}\left(\mathcal{W}_{t}\right)\}$ with that of accuracy during the learning dynamics, varying with the level of guessing bias at initialization (IGB). Particularly, $\gamma(\mathcal{A},\psi(\chi))\in\mathbb{R}^{+}$ (colormap on the right) provides a measure of the level of IGB: the higher the value of $\gamma(\mathcal{A},\psi(\chi))$, the higher the level of IGB (see Sec. 4 for more details). The curves show a consistent pattern with the diagram on the left. They also demonstrate that the time for IGB absorption increases with the level of IGB itself. The simulations were conducted on an MLP-mixer using a binary dataset (dogs vs cats from CIFAR) as input; more details on the setting and additional experiments with more architectures/datasets are provided in App. I.2.
@@
We start with an informal definition of the phenomenon of IGB. Consider a dataset $\chi$ and an architecture $\mathcal{A}$, parameterized by a set of weights. Specifically, we denote the configuration of the weight set at time $t$ as $\mathcal{W}_{t}$, with $\mathcal{W}\equiv\mathcal{W}_{0}$. An observable that characterizes the IGB phenomenon is the fraction of points classified, by the untrained model, as class $c$, and denoted by $f_{c}\left(\mathcal{W}\right)$.
@@
We say there is an Initial Guessing Bias if there is an imbalance in the initial fractions of points, captured by $f_{c}\left(\mathcal{W}\right)$, assigned to a class.
@@
As we will show in this paper, IGB is a pervasive phenomenon emerging in untrained models. Therefore, a natural question arises: what are the consequences of IGB in terms of training and generalization? Fully addressing this question and determining whether IGB has beneficial or detrimental effects on network training is complex as the answer might vary with the problem (data, architecture, optimizer, choices of hyperparameter, and so on might all have confounding effects). A comprehensive study of this kind is beyond the scope of this work. However, we can start with a simple observation; a discrepancy between the fractions of guesses, $f_{c}\left(\mathcal{W}_{t}\right)$}, and the actual class proportions in the dataset, sets a limit on the achievable performance. For instance, considering a balanced dataset with $N_{C}$ classes, we can derive the following upper bound on the accuracy (Fig. 2 (left)):
$\text{Accuracy}(t)\leq 1-\left(\max_{c}\{f_{c}\left(\mathcal{W}_{t}\right)\}-\frac{1}{N_{C}}\right).$ (1)
@@ 4 Emergence of IGB: a first insight into the phenomenon
Multi-Layer Perceptron (MLP) submodules are widely integrated into various architectures (He et al., 2016; Vaswani et al., 2017; Dosovitskiy et al., 2020). Moreover, recent advances, such as the MLP-mixer (Tolstikhin et al., 2021), highlight MLPs’ capability to achieve competitive performance, suggesting their untapped potential, especially in large dataset applications. Our theoretical analysis centers on understanding the intricacies of MLP architectures. Formally, we consider an MLP as a set of $D$ inputs $\{\boldsymbol{\xi}^{(a)}\}_{a=1}^{D}$ that propagate through a set of $L$ hidden layers, until they reach the output layer, composed of a set of two nodes (one for each class), $\left\{O^{(c)}\right\}_{c=0,1}\,$ (see App. A for more details). Each input is classified by selecting the class $c$ with the largest output value.
@@
In our analysis, the random variables (r.v.s) we consider are essentially functions of two independent sources of randomness: the dataset, $\chi$, and the set of initialized weights, $\mathcal{W}$. We denote the cumulative distribution function (c.d.f.) of a r.v. $X$ as $F_{X}\left({x}\right)$ and its probability density function (p.d.f.) as $p_{X}\left({x}\right)$. For variables dependent on both sources of randomness, we specify the active source in the notation. For instance, given a function of two independent sets of random variables, $X(\mathcal{W},\chi)$:
$F_{X}^{(\chi)}\left({x}\right)=\mathbb{P}\left(X<x|\mathcal{W}\right)\,,\,\,p_{X}^{(\chi)}\left({x^{\prime}}\right)=\frac{d}{dx}F_{X}^{(\chi)}\left({x}\right)\Bigr|_{\begin{subarray}{c}x=x^{\prime}\end{subarray}},$ (2)
@@
We study how untrained neural networks assign data points to different classes before they start learning. In the absence of explicit bias weights, our intuition might suggest an even partitioning between classes. Instead, we demonstrate that the model design choices can induce an imbalance effect, resulting in a large fraction of data points being classified as a single class. The key quantity in our analysis is the distribution (over $\mathcal{W}$) of the fraction $f_{c}\left(\mathcal{W}\right)$ of datapoints classified as class $c\in\{0,1\}$. We have $f_{c}=1/2$ in the absence of IGB, and a different value otherwise.
Since our model’s guess is assigned to the class with the largest output value, $O^{(c)}$, in the limit of infinite datapoints, the Law of large numbers implies that
$\lim_{D\rightarrow\infty}f_{c}\left(\mathcal{W}\right)=\mathbb{P}\left(O^{(c)}>O^{(1-c)}\mid\mathcal{W}\right)=$ (3)
$\int_{0}^{\infty}p_{\Delta_{O^{(c)}}}^{(\chi)}\left({x}\right)\;dx\,,\,\,\text{with}\,\,\Delta_{O^{(c)}}\equiv O^{(c)}-O^{(1-c)}.$
@@
$\lim_{D\to\infty}f_{0}\left(\mathcal{W}\right)=\mathbb{P}\left(\Delta_{O}>0\mid\mathcal{W}\right)=\int_{0}^{\infty}p_{\Delta_{O}}^{(\chi)}\left({x}\right)\;dx\,$, (4)
with $\Delta_{O}\equiv O^{(0)}-O^{(1)}$. For the sake of compactness, we define $\mu_{c}\equiv\left\langle{O^{(c)}}\right\rangle$ and $\Delta_{\mu}\equiv\mu_{0}-\mu_{1}$ as the difference between the distributions mean values. While $f_{c}$ is a convenient and interpretable metric related to performance, our analysis can incorporate alternative measures that, for instance, allow for the analysis of the confidence level in the network’s assignments, as shown in App. E.2.
Eq. (4) connects $f_{0}$ (i.e. the observable we are interested in) to the nodes variables $\left\{O^{(c)}\right\}$ (i.e. the set of variables we analyze through our investigation). The value of $f_{0}$ depends on how often the output related to class 0 has a higher value of the output than that of class $1$. More precisely, Eq. (4) shows that $f_{0}$ is determined by comparing the output distributions $p_{O^{(c)}}^{(\chi)}\left({x}\right)$ related to each class $c$. We illustrate this in Fig. 3. In the example of the figure, the output distribution related to class 0 is centered around higher values than that of class 1. Therefore, we will have $f_{0}>f_{1}$, indicating the presence of IGB. In App. D.1 we show that, for MLPs, $p_{O^{(c)}}^{(\chi)}\left({x}\right)$ , is asymptotically a Gaussian whose center, $\mu_{c}$ , is itself a r.v., which is drawn from a Normal distribution $p_{\mu_{c}}\left({x}\right)$ that has a wide support:
@@
Consider a Gaussian distributed dataset processed through an MLP with $L$ hidden layers and weights initialized according to the Kaiming normal scheme (with null bias weights). In the limit of infinite width, the distribution of an output node $O^{(c)}$, at initialization, converges to: $p_{O^{(c)}}^{(\chi)}\left({x}\right)\xrightarrow{|\mathcal{W}|\rightarrow\infty}\mathcal{N}\left(x;\mu_{c},\textrm{Var}_{\chi}\left({O^{(c)}}\right)\right),$ (5) where $|\mathcal{W}|$ indicates the cardinality of the set $\mathcal{W}$ and, for compactness, we denoted by $|\mathcal{W}|\rightarrow\infty$ the limit where the number of neurons, $N_{l}$, of each hidden layer $l\in\{0,\dots,L\}$ tends to infinity.
Moreover, while the variance of the distribution, $\textrm{Var}_{\chi}\left({O^{(c)}}\right)$, converges with high probability (w.h.p.) to a deterministic value, the center of this distribution, $\mu_{c}$, is itself a r.v., varying from node to node and converging in distribution to:
$p_{\mu_{c}}\left({x}\right)\xrightarrow{|\mathcal{W}|\rightarrow\infty}\mathcal{N}\left(x;0,\textrm{Var}_{\mathcal{W}}\left({\mu_{c}}\right)\right).$ (6)
@@
In fact, the study of IGB can be conceptually summarized as a comparison between the fluctuations of $p_{\mu_{c}}\left({x}\right)$ , which defines the distance between the Gaussians in Fig. 3, and those of $p_{O^{(c)}}^{(\chi)}\left({x}\right)$ , which define how wide each of these Gaussians is. We now illustrate the two extreme cases to underline our point. Starting from Eq. (3), we will discuss how the integral on the r.h.s. varies in these two scenarios:
@@
    If the fluctuations of $p_{\mu_{c}}\left({x}\right)$ are much smaller than the ones of $p_{O^{(c)}}^{(\chi)}\left({x}\right)$ , we will have two Gaussian r.v.s centered almost on the same point, therefore, $\mathbb{P}\left(O^{(0)}>O^{(1)}\mid\mathcal{W}\right)=\mathbb{P}\left(\Delta_{O}>0\mid\mathcal{W}\right)\simeq 1/2$ .
    Indeed, the difference between two Gaussian r.v.s is itself a Gaussian r.v., centered at the difference between the mean values of the original distributions, $\Delta_{\mu}$ . If the fluctuations of $\Delta_{\mu}$ are much smaller than those of $\Delta_{O}$, we will typically have that $p_{O^{(c)}}^{(\chi)}\left({x}\right)$ is a symmetric distribution centered very close to the origin. Therefore $\mathbb{P}\left(\Delta_{O}>0\mid\mathcal{W}\right)\simeq 1/2$ i.e., from Eq. (3), the fraction of points assigned to both classes is equal to 1/2.
@@
    If, instead, the scale of $p_{\mu_{c}}\left({x}\right)$ fluctuations is much bigger than that of $p_{O^{(c)}}^{(\chi)}\left({x}\right)$ we will typically fall in the opposing scenario where the two Gaussian distributions, $p_{O^{(c)}}^{(\chi)}\left({x}\right)$ , are well separated. We can assume, without loss of generality, that $\mu_{0}>\mu_{1}$ . In this case we will have $\mathbb{P}\left(O^{(0)}>O^{(1)}\mid\mathcal{W}\right)=\mathbb{P}\left(\Delta_{O}>0\mid\mathcal{W}\right)\simeq 1$ .
@@
Given an architecture $\mathcal{A}$ and a preprocessed dataset $\psi(\chi)$, we have an absence of IGB if and only if, w.h.p., $\lim_{D\rightarrow\infty}f_{c}\left(\mathcal{W}\right)=\frac{1}{N_{C}},\;\forall c\in\{0,\dots,N_{C}-1\}.$ (7) We instead have a presence of IGB if, in the limit of infinite data points ($D\rightarrow\infty$), we observe a disproportion between the values $\{f_{c}\}$ for different classes.
Note that, when $D$ is finite, there can be finite-size fluctuations which move $f_{c}\left(\mathcal{W}\right)$ from its asymptotic value.
@@
$\gamma(\mathcal{A},\psi(\chi))\equiv\frac{\textrm{Var}_{\mathcal{W}}\left({\mu_{c}}\right)}{\textrm{Var}_{\chi}\left({O^{(c)}}\right)}\,.$ (8)
@@
$\gamma(\mathcal{A},\psi(\chi))=\infty\Longrightarrow p_{f_{0}}\left({x}\right)=\tfrac{1}{2}\delta\left(x\right)+\tfrac{1}{2}\delta\left(x-1\right),$ (10)
indicating that in each experiment, the dataset is completely classified as either belonging to class 0 or class $1$.

### v5

@@ 1 Introduction
Figure 1: Initial Guessing Bias (IGB). Consider a task where we classify a binary dataset using an untrained network. Does it assign half of the examples to each class, or does it privilege one class? The answer depends on the model design. In the top-left, we classify a binary dataset with an untrained network without IGB. This model will generally assign half of the examples to each class (histogram on the top-center). In the bottom-left, we classify the same dataset using an untrained network with IGB. In this case, most of the guesses will usually go to one of the two classes (histogram on the bottom-center). As an example, we take the dog/cat classes (label 0 / label 11) from CIFAR10, and pass them through an untrained CNN with 2 layers, each followed by pooling. The non-IGB model uses $\tanh$ activations and average pooling, the IGB model uses ReLU and max pooling. We show in the center-right the distribution over different initializations, $f_{G_{0}}\left({g_{0}}\right)$, of the fraction $G_{0}$ of times that each model guessed dog (equivalently, $G_{1}=1-G_{0}$ indicates the fraction of images guessed as cat). While for the non-IGB models, $G_{0}$ is most often 50%, with IGB it most often is either 0% or 100%.
@@ 3 Preliminaries
Figure 2: Left: IGB and Performance Bound: Diagram of the accessible performance range conditioned on the behavior of $\max_{c}\{G_{c}\left(\mathcal{W}_{t}\right)\}$ in a balanced binary dataset, in accordance with Eq. (1). Right: Comparison of the trend of $\max_{c}\{G_{c}\left(\mathcal{W}_{t}\right)\}$ with that of accuracy during the learning dynamics, varying with the level of guessing bias at initialization (IGB). Particularly, $\gamma(\mathcal{A},\psi(\chi))\in\mathbb{R}^{+}$ (colormap on the right) provides a measure of the level of IGB: the higher the value of $\gamma(\mathcal{A},\psi(\chi))$, the higher the level of IGB (see Sec. 4 for more details). The curves show a consistent pattern with the diagram on the left. They also demonstrate that the time for IGB absorption increases with the level of IGB itself. The simulations were conducted on an MLP-mixer using a binary dataset (dogs vs cats from CIFAR) as input; more details on the setting and additional experiments with more architectures/datasets are provided in App. I.2.
@@
We start with an informal definition of the phenomenon of IGB. Consider a dataset $\chi$ and an architecture $\mathcal{A}$, parameterized by a set of weights. Specifically, we denote the configuration of the weight set at time $t$ as $\mathcal{W}_{t}$, with $\mathcal{W}\equiv\mathcal{W}_{0}$. An observable that characterizes the IGB phenomenon is the fraction of points classified, by the untrained model, as class $c$, and denoted by $G_{c}\left(\mathcal{W}\right)$.
@@
We say there is an Initial Guessing Bias if there is an imbalance in the initial fractions of points, captured by $G_{c}\left(\mathcal{W}\right)$, assigned to a class.
@@
As we will show in this paper, IGB is a pervasive phenomenon emerging in untrained models. Therefore, a natural question arises: what are the consequences of IGB in terms of training and generalization? Fully addressing this question and determining whether IGB has beneficial or detrimental effects on network training is complex as the answer might vary with the problem (data, architecture, optimizer, choices of hyperparameter, and so on might all have confounding effects). A comprehensive study of this kind is beyond the scope of this work. However, we can start with a simple observation; a discrepancy between the fractions of guesses, $G_{c}\left(\mathcal{W}_{t}\right)$}, and the actual class proportions in the dataset, sets a limit on the achievable performance. For instance, considering a balanced dataset with $N_{C}$ classes, we can derive the following upper bound on the accuracy (Fig. 2 (left)):
$\text{Accuracy}(t)\leq 1-\left(\max_{c}\{G_{c}\left(\mathcal{W}_{t}\right)\}-\frac{1}{N_{C}}\right).$ (1)
@@ 4 Emergence of IGB: a first insight into the phenomenon
Multi-Layer Perceptron (MLP) submodules are widely integrated into various architectures (He et al., 2016; Vaswani et al., 2017; Dosovitskiy et al., 2020). Moreover, recent advances, such as the MLP-mixer (Tolstikhin et al., 2021), highlight MLPs’ capability to achieve competitive performance, suggesting their untapped potential, especially in large dataset applications. Our theoretical analysis centers on understanding the intricacies of MLP architectures. Formally, we consider an MLP as a set of $D$ inputs $\{\boldsymbol{\xi}^{(a)}\}_{a=1}^{D}$ that propagate through a set of $L$ hidden layers, until they reach the output layer, composed of a set of two nodes (one for each class), $\left\{O_{c}\right\}_{c=0,1}\,$ (see App. A for more details). Each input is classified by selecting the class $c$ with the largest output value.
@@
In our analysis, the random variables (r.v.s) we consider are essentially functions of two independent sources of randomness: the dataset, $\chi$, and the set of initialized weights, $\mathcal{W}$. We denote the cumulative distribution function (c.d.f.) of a r.v. $X$ as $F_{X}\left({x}\right)$ and its probability density function (p.d.f.) as $f_{X}\left({x}\right)$. For variables dependent on both sources of randomness, we specify the active source in the notation. For instance, given a function of two independent sets of random variables, $X(\mathcal{W},\chi)$:
$F_{X}^{(\chi)}\left({x}\right)=\mathbb{P}\left(X<x|\mathcal{W}\right)\,,\,\,f_{X}^{(\chi)}\left({x^{\prime}}\right)=\frac{d}{dx}F_{X}^{(\chi)}\left({x}\right)\Bigr|_{\begin{subarray}{c}x=x^{\prime}\end{subarray}},$ (2)
@@
We study how untrained neural networks assign data points to different classes before they start learning. In the absence of explicit bias weights, our intuition might suggest an even partitioning between classes. Instead, we demonstrate that the model design choices can induce an imbalance effect, resulting in a large fraction of data points being classified as a single class. The key quantity in our analysis is the distribution (over $\mathcal{W}$) of the fraction $G_{c}\left(\mathcal{W}\right)$ of datapoints classified as class $c\in\{0,1\}$. We have $G_{c}=1/2$ in the absence of IGB, and a different value otherwise.
Since our model’s guess is assigned to the class with the largest output value, $O_{c}$, in the limit of infinite datapoints, the Law of large numbers implies that
$\lim_{D\rightarrow\infty}G_{c}\left(\mathcal{W}\right)=\mathbb{P}\left(O_{c}>O_{1-c}\mid\mathcal{W}\right)=$ (3)
$\int_{0}^{\infty}f_{\Delta_{O_{c}}}^{(\chi)}\left({\delta_{o}}\right)\;d\delta_{o}\,,\,\,\text{with}\,\,\Delta_{O_{c}}\equiv O_{c}-O_{1-c}.$
@@
$\lim_{D\to\infty}G_{0}\left(\mathcal{W}\right)=\mathbb{P}\left(\Delta_{O}>0\mid\mathcal{W}\right)=\int_{0}^{\infty}f_{\Delta_{O}}^{(\chi)}\left({x}\right)\;dx\,$, (4)
with $\Delta_{O}\equiv O_{0}-O_{1}$. For the sake of compactness, we define $\mu_{c}\equiv\left\langle{O_{c}}\right\rangle$ and $\Delta_{\mu}\equiv\mu_{0}-\mu_{1}$ as the difference between the distributions mean values. While $G_{c}$ is a convenient and interpretable metric related to performance, our analysis can incorporate alternative measures that, for instance, allow for the analysis of the confidence level in the network’s assignments, as shown in App. E.2.
Eq. (4) connects $G_{0}$ (i.e. the observable we are interested in) to the nodes variables $\left\{O_{c}\right\}$ (i.e. the set of variables we analyze through our investigation). The value of $G_{0}$ depends on how often the output related to class 0 has a higher value of the output than that of class 11. More precisely, Eq. (4) shows that $G_{0}$ is determined by comparing the output distributions $f_{O_{c}}^{(\chi)}\left({o}\right)$ related to each class $c$. We illustrate this in Fig. 3. In the example of the figure, the output distribution related to class 0 is centered around higher values than that of class 1. Therefore, we will have $G_{0}>G_{1}$, indicating the presence of IGB. In App. D.1 we show that, for MLPs, $f_{O_{c}}^{(\chi)}\left({o}\right)$ , is asymptotically a Gaussian whose center, $\mu_{c}$ , is itself a r.v., which is drawn from a Normal distribution $f_{\mu_{c}}\left({m}\right)$ that has a wide support:
@@
Consider a Gaussian distributed dataset processed through an MLP with $L$ hidden layers and weights initialized according to the Kaiming normal scheme (with null bias weights). In the limit of infinite width, the distribution of an output node $O_{c}$, at initialization, converges to: $f_{O_{c}}^{(\chi)}\left({o}\right)\xrightarrow{|\mathcal{W}|\rightarrow\infty}\mathcal{N}\left(o;\mu_{c},\textrm{Var}_{\chi}\left({O_{c}}\right)\right),$ (5) where $|\mathcal{W}|$ indicates the cardinality of the set $\mathcal{W}$ and, for compactness, we denoted by $|\mathcal{W}|\rightarrow\infty$ the limit where the number of neurons, $N_{l}$, of each hidden layer $l\in\{0,\dots,L\}$ tends to infinity.
Moreover, while the variance of the distribution, $\textrm{Var}_{\chi}\left({O_{c}}\right)$, converges with high probability (w.h.p.) to a deterministic value, the center of this distribution, $\mu_{c}$, is itself a r.v., varying from node to node and converging in distribution to:
$f_{\mu_{c}}\left({m}\right)\xrightarrow{|\mathcal{W}|\rightarrow\infty}\mathcal{N}\left(m;0,\textrm{Var}_{\mathcal{W}}\left({\mu_{c}}\right)\right).$ (6)
@@
In fact, the study of IGB can be conceptually summarized as a comparison between the fluctuations of $f_{\mu_{c}}\left({m}\right)$ , which defines the distance between the Gaussians in Fig. 3, and those of $f_{O_{c}}^{(\chi)}\left({o}\right)$ , which define how wide each of these Gaussians is. We now illustrate the two extreme cases to underline our point. Starting from Eq. (3), we will discuss how the integral on the r.h.s. varies in these two scenarios:
@@
    If the fluctuations of $f_{\mu_{c}}\left({m}\right)$ are much smaller than the ones of $f_{O_{c}}^{(\chi)}\left({o}\right)$ , we will have two Gaussian r.v.s centered almost on the same point, therefore, $\mathbb{P}\left(O_{0}>O_{1}\mid\mathcal{W}\right)=\mathbb{P}\left(\Delta_{O}>0\mid\mathcal{W}\right)\simeq 1/2$ .
    Indeed, the difference between two Gaussian r.v.s is itself a Gaussian r.v., centered at the difference between the mean values of the original distributions, $\Delta_{\mu}$ . If the fluctuations of $\Delta_{\mu}$ are much smaller than those of $\Delta_{O}$, we will typically have that $f_{O_{c}}^{(\chi)}\left({o}\right)$ is a symmetric distribution centered very close to the origin. Therefore $\mathbb{P}\left(\Delta_{O}>0\mid\mathcal{W}\right)\simeq 1/2$ i.e., from Eq. (3), the fraction of points assigned to both classes is equal to 1/2.
@@
    If, instead, the scale of $f_{\mu_{c}}\left({m}\right)$ fluctuations is much bigger than that of $f_{O_{c}}^{(\chi)}\left({o}\right)$ we will typically fall in the opposing scenario where the two Gaussian distributions, $f_{O_{c}}^{(\chi)}\left({o}\right)$ , are well separated. We can assume, without loss of generality, that $m_{0}>m_{1}$ . In this case we will have $\mathbb{P}\left(O_{0}>O_{1}\mid\mathcal{W}\right)=\mathbb{P}\left(\Delta_{O}>0\mid\mathcal{W}\right)\simeq 1$ .
@@
Given an architecture $\mathcal{A}$ and a preprocessed dataset $\psi(\chi)$, we have an absence of IGB if and only if, w.h.p., $\lim_{D\rightarrow\infty}G_{c}\left(\mathcal{W}\right)=\frac{1}{N_{C}},\;\forall c\in\{0,\dots,N_{C}-1\}.$ (7) We instead have a presence of IGB if, in the limit of infinite data points ($D\rightarrow\infty$), we observe a disproportion between the values $\{G_{c}\}$ for different classes.
Note that, when $D$ is finite, there can be finite-size fluctuations which move $G_{c}\left(\mathcal{W}\right)$ from its asymptotic value.
@@
$\gamma(\mathcal{A},\psi(\chi))\equiv\frac{\textrm{Var}_{\mathcal{W}}\left({\mu_{c}}\right)}{\textrm{Var}_{\chi}\left({O_{c}}\right)}\,.$ (8)
@@
$\gamma(\mathcal{A},\psi(\chi))=\infty\Longrightarrow f_{G_{0}}\left({g_{0}}\right)=\tfrac{1}{2}\delta\left(g_{0}\right)+\tfrac{1}{2}\delta\left(g_{0}-1\right),$ (10)
indicating that in each experiment, the dataset is completely classified as either belonging to class 0 or class 11.